# Human2Robot: Learning Robot Actions from Paired Human-Robot Videos

### v1

@@ 4.2 Training
Our approach consists of two training stages. In the first stage, we focus on image generation by taking the first frame of $o_{r}^{t}$, the first frame of human $o_{h}^{t}$ and future frame of human $o_{h}^{t+i}$ as inputs to predict the future frame of robot $\hat{o}_{r}^{t+i}$. In this stage, we train the S-UNet, the Behavior Extractors, and the ST-UNet without the temporal layers. The spatial layers in both the ST-UNet and S-UNet are instantiated using the powerful open-source Stable Diffusion model (SD) [50], while the Behavior Extractors are initialized with Gaussian weights, except for the final projection layer, which utilizes zero initialized convolution.
In the second stage, we concentrate the training effort on the temporal layer. In this stage, we take a 30-frame segment of human video $o_{h}^{t:t+n}$ and the first frame $o_{r}^{t}$ of the robot video $o_{r}^{t:t+n}$ as inputs to predict the future robot observation $\hat{o}_{r}^{t:t+n}$ and action $a_{t:t+n}$. The first frame of the robot video is fed into Behavior Extractor 1, while the entire video segment is passed into Behavior Extractor 2, yielding features $x_{1}^{t}$ and $x_{2}^{t:t+n}$, respectively. Then, we replicate $x_{1}^{t}$ by n times and add it to $x_{2}^{t:t+n}$, which is subsequently added to noise latent and fed to our ST-UNet. The weights of the VAE Encoder and Decoder, as well as the CLIP image encoder are frozen all the time. During our training process, we find that dense sampling (30fps) results in jittering in generated videos. Therefore, we reduce the frame rate to one-fourth of the original, namely cutting the number of frames by three-quarters. Optimization Objectives. We train Human2Robot with both generation and action prediction objectives, aiming for the model to achieve high-quality video generation and accurate action prediction simultaneously. The goal of generation for the model is to predict the applied noise, so the objective of the optimization process is:
@@
the parameters of the formula represent the latent state $z_{ts}$ at time step $ts$, which is obtained from Spatial-Temporal UNet. The conditioning variable $c$ includes the first frame $o_{r}^{t}$ of the robot and a human video $o_{h}^{t:t+n}$. $\epsilon$ denotes the added noise, and $\epsilon_{\theta}(z_{ts},c,ts)$ represents the model’s predicted noise.
@@
Method Seen Task Changed Task Unseen Task
Task-1 Task-2 Task-3 Task-4 Task-5 Task-6 Task-7 Task-8
Demonstration Videos Retrieved by KNN
ACT* (KNN) [71] 75 55 0 80 30 - - -
Human2Robot (KNN) 60 85 80 70 60 - - -
Manual Demonstration Videos
ACT* [71] 70 0 0 85 20 45 0 0
Human2Robot 85 95 90 80 75 75 95 85
Table 1: Quantitative evaluation results of KNN and manual guidance. KNN-based methods are not able to complete different tasks with the same initial scene (Task-6) and unseen tasks (Task-7 Task-8). The task description is as follows: Seen Tasks: (1) Pinch cube: pinch the cube and place it on the plate. (2) Push box: push the box from left to right. (3) Push plate: push the plate from top to bottom. Changed Tasks: (4) Pinch a pen on the plate: pinch a pen in a plate and place it on the plate left to it (unseen background, with no plate below the pen in the training set). (5) Pinch a ping pong ball: pinch the ping pong ball (unseen object) and place it on the plate. (6) Pinch a cube to the designated plate (bottom/top): pinch a cube and place it on the bottom/top plate (different trajectories). Unseen Tasks: (7) Pinch a banana from bottom to top: with a plate in the scene (similar to the tasks in the training set), pinch the banana (unseen object) from bottom to top in the image (unseen trajectory). (8) Rotate once before pushing the box: Turn the hand 360° counterclockwise (unseen trajectory) and push the box.
@@ 4.3 Inference
Using KNN significantly simplifies the process for known tasks. However, for tasks involving different objectives within the same setting—such as Task 6 in Table 1, as well as for unseen tasks, explicit human video guidance is required to accomplish these tasks effectively. Temporal Addition of Actions. To mitigate trajectory jittering during sequential prediction, inspired by the approach used in ACT [71], we derive the action not from a single prediction but from a weighted average of multiple prediction steps. In particular, we predict a series of actions $a_{t:t+n}$ at time t, which are then combined with the previously predicted actions to compute the aggregated actions $A_{t:t+n}^{t}$. This represents the set of actions predicted at time $t$, to be executed from time $t$ to $t+n$, and can be computed as follows:
@@
where $\gamma$ is a hyperparameter, with a larger $\gamma$ indicating a stronger emphasis on the current prediction, and $m$ ($m\leq n$) refers to the length of the action executed after each prediction. Specifically, when $m=n$, the inference degenerates into sequential prediction. Inference Acceleration. During inference, the diffusion model typically goes through $T$ denoising steps to generate the output. We explore the effectiveness of using features from different denoising steps in Section 5.3, and find that features obtained from the first half of the denoising stages lead to better results. Therefore, to balance speed and quality, we halt the denoising process after obtaining features from the $\frac{T}{4}$ step, directly predict and execute the action, without waiting for the entire denoising process to complete. It is worth noting that if we aim to achieve the fastest possible speed, we can directly use the one-step denoising features for prediction, which allows us to predict the next 30 steps of actions within 0.6 seconds.
@@ 5.1 Experimental Setups
Task Definition. We conduct testing across eight experiments, as depicted in Table 1. These experiments are categorized based on their familiarity and complexity into three types: Seen Tasks are present during the training phase but feature slight variations, such as changes in object positions during testing. Changed Tasks are designed to test the generalization capabilities of the model, as it requires the model to understand human hand movements and transfer actions to different types of objects. Unseen Tasks introduce completely new trajectories. Previous methods [71, 39] struggle to achieve this goal because completing a task with an unknown trajectory is highly challenging.
Baseline Construction. We compare Human2Robot with the following baselines. All baselines and Human2Robot utilize the same H&R dataset for training across all tasks for a fair comparison.
  • ACT* [71]: For a fair comparison, we modify ACT to a video conditioned policy, and we name it ACT*. Details can be found in the Section D.
  • XSkill [68]: XSkill is a human-video-conditioned policy managed to bridge the embodiment gap between human and robot.
pinch cube pull plate pinch and pull (unseen)
XSkill [68] 40 60 0
Human2Robot 85 90 60
Table 2: Quantitative results comparison between XSkill and Human2Robot. We compare the basic ability on seen tasks and the generalization ability on unseen tasks (combination of seen tasks).
Evaluation Protocols. For the evaluation of the KNN-based methods, we evaluate models using seen tasks by randomly positioning objects within the defined area during testing. This evaluates the KNN’s effectiveness in identifying and successfully completing the known tasks. For testing scenarios in which the hand demonstration videos are manually given, we collect 10 new data samples, ensuring variation in position, trajectory, and lighting conditions. Additionally, we collect 20 demonstrations for each result to calculate the success rate.
Figure 4: Qualitative visualization of generated result and robot action result.
@@ 5.2 Main Results
We discuss the capabilities of the model in generating realistic robot videos and robot manipulation experiments. The evaluation tasks vary from seen to changed and unseen tasks, allowing for a comprehensive study of our method’s imitative ability and generalization.
Action Prediction with KNN. We test the model in a setting where no paired demonstrations are provided. To gain the proper guidance signals, we employ KNN, which utilizes features extracted from DINOv2 [44] and CLIP [46] for searching the high-quality and relevant episodes as hand demonstration videos. As shown in Table 1, Human2Robot performs well across multiple seen tasks, achieving 85% and 80% on tasks 2 and 3, respectively. In comparison, the ACT* model fails to complete these tasks since ACT is not designed for multi-task learning. In particular, due to the prevalence of grasping tasks in our dataset, ACT performs well on grasping tasks (Task-1, Task-4), but poorly on pushing tasks (Task-2, Task-3).
Besides, ACT places more emphasis on the initial scene information than human videos, consistent with its design purpose, which results in it being able to execute tasks based on the initial scene even when the KNN selects an incorrect human hand video. In some cases (Task-1, Task-2, Task-5), its success rate even exceeds that of the ground truth. However, Our model is more reliant on the human hand video, which means that when the KNN selects a less suitable demonstration, the model is likely to fail, leading to poor performance on Task-1.
Moreover, in the Changed Task setting, Human2Robot performs a comparable or better successful rate than ACT* model, achieving 70% on Task-4 and 60% on Task-5, while ACT tends to fail due to the unseen initial frame. Such competitive results demonstrate the high efficiency and accuracy of Human2Robot in action prediction with episodes retrieved by KNN.
Action Prediction with Human Demonstration Videos. We also evaluate performance in scenarios where paired human demonstration videos are provided. As shown in Table 2, Human2Robot outperforms XSkill [68] in both seen and unseen tasks, highlighting the superiority of our model. Besides, as shown in Table 1, Human2Robot significantly benefits from using these GT demonstrations instead of relying solely on KNN, with an improvement in success rates of over 10%. This enhancement is largely attributed to the test set containing some positions not seen during training, which poses a challenge for KNN in finding appropriate demonstrations. Additionally, Human2Robot consistently outperforms ACT* across all seen tasks, unseen tasks, and most changed tasks. Notably, in the Unseen Task category which includes challenging actions such as moving a banana from the bottom to the top of the frame and rotating in a circle before pushing a box—Human2Robot achieves impressive success rates of 95% and 85%, respectively, whereas ACT* struggles to generalize in such condition. These results highlight the strong generalization capabilities of our model, showcasing its ability to perform tasks with significantly greater accuracy than ACT*.
Moreover, on more complex 6DOF task with significant background changes, Human2Robot can also achieve an accuracy of 80%, as shown in Figure 5. For the handwriting task, having one-shot generalization ability for unseen tasks is crucial, as it is impossible to cover all possible characters in the training data. However, to the best of our knowledge, no existing model has achieved this level of generalization—except for Human2Robot, as shown in Figure 6.
Realistic Robot Video Generation. Besides predicting actions, Human2Robot can visually simulate future dynamics of robot movements, requiring precise predictions of the motions of the robot arm and its interactions with objects. As depicted in Figure 4, Human2Robot effectively generates realistic robot arm videos based on the initial state and human demonstrations. Impressively, these videos not only mirror human hand movements but are also consistently accurate in representing real-world robotic movements and interactions. These results highlight the strong imaginative capability of our generative model, demonstrating its ability to bridge visual discrepancies between human and robot arms. It also accurately maps the relationship between human hand movements, robot actions, and object interactions.
Figure 5: Qualitative visualization of upright brush task. A full 6-DOF task with significant background changes and Human2Robot achieves an 80% success rate on this task.
@@ 5.3 Discussion
Figure 6: Handwriting task. Trained on random movements data and one-shot to unseen character writing in brand new background.
Generated Video Usage and Evaluation. We further explore the capabilities of Human2Robot by generating demonstrations for a robot-video-conditioned policy using human demonstration videos. In this case, Human2Robot serves to expand the input sources, allowing the use of human videos instead of robot videos, which is more convenient. Specifically, we choose ACT* as the test video-conditioned policy. We use the robot-video-conditioned model, instead of the human-video-conditioned model in Table 1. However, during inference, condition it with robot demonstrations (w Robot GT) and robot demonstrations generated by Human2Robot (w Generated) from human video. As shown in Table 3, the performance of our generated video is competitive with that of the robot video.
Moreover, to further evaluate the quality of our generated video, we treat the above method as a way to assess video quality, measuring the difference between generated and real videos. We regard the performance of robot video as an upper bound. The performance when using generated videos during inference serves as a proxy for evaluating visual quality. As shown in the Table 3 below, replacing GT videos with our generated videos achieves a similar score, demonstrating the realism of our video prediction.
Task1 Task2 Task3 Task4 Task6 Task8 (unseen)
ACT* (w Robot GT) 80 90 0 80 60 0
ACT* (w Generated) 60 80 0 80 40 0
Human2Robot 85 95 90 80 75 85
Table 3: Performance comparison between robot demonstration and robot demonstration generated by Human2Robot from human demonstration.
The Choice of Upsampling Layers for Feature Extraction. As shown in Table 4, we examine the impact of extracting features from different UNet upsample layers for predicting actions via the MLP. Using the second up-sampling block (Up Index 2) achieves a 90% success rate on seen tasks, while the first up-sampling block (Up Index 1) reaches a 95% success rate on unseen tasks. The results indicate that the choice of upsampling layer has minimal effect on outcomes.
@@
Studying the Choice of Denoising Step. As shown in Tab 5, During inference, we fix the upsampling layer to extract features from different denoising steps. We find that, due to the MLP being trained with random noise, features from just one denoising step already allow the model to perform well, achieving 80% accuracy on seen tasks and 90% on unseen tasks, with an inference time of 0.6 seconds for 30 step trajectory prediction. Features after 7 denoising steps (1/4 of the total steps) yield the highest success rate. However, the results are worse at 15 (1/2 total steps) and 22 (3/4 total steps), possibly due to the following reason: in early denoising stages, as shown in Figure 7, the feature focuses more on position information which enhances the success rate. However, in later denoising stages, the model focuses more on reconstructing detailed visual information, which increases action variance, making the movements jerky of the robot and ultimately reducing the success rate.
Figure 7: Visual result of one-step DDIM.
@@ 6 Conclusion
In this paper, we introduced a data collection pipeline for perfectly paired human-robot data and used it to create the H&R dataset. Then we presented Human2Robot, a novel end-to-end diffusion framework that simultaneously predicted robot video and actions given a hand video. Next, we leveraged a KNN method to perform tasks without explicitly requiring human hand videos as input. Human2Robot performed excellently not only on seen tasks but also on changed and unseen tasks and all the tasks can be executed in real time.

### v2

@@ 4.2 Training
Our approach consists of two training stages. In the first stage, we focus on image generation by taking the first frame of $o^{r}_{t}$, the first frame of human $o^{h}_{t}$ and future frame of human $o^{h}_{t+i}$ as inputs to predict the future frame of robot $o^{r}_{t+i}$. In this stage, we train the S-UNet, the Behavior Extractors, and the ST-UNet without the temporal layers. The spatial layers in both the ST-UNet and S-UNet are instantiated using the powerful open-source Stable Diffusion model (SD)  [48], while the Behavior Extractors are initialized with Gaussian weights, except for the final projection layer, which utilizes zero initialized convolution.
In the second stage, we concentrate the training effort on the temporal layer. In this stage, we take a 30-frame segment of human video $o^{h}_{t:t+n}$ and the first frame $o^{r}_{t}$ of the robot video $o^{r}_{t:t+n}$ as inputs to predict the future robot observation $o^{r}_{t:t+n}$ and action $a_{t:t+n}$. The first frame of the robot video is fed into Behavior Extractor 1, while the entire video segment is passed into Behavior Extractor 2, yielding features $x_{1}^{t}$ and $x_{2}^{t:t+n}$, respectively. Then, we replicate $x_{1}^{t}$ by n times and add it to $x_{2}^{t:t+n}$, which is subsequently added to noise latent and fed to our ST-UNet. The weights of the VAE Encoder and Decoder, as well as the CLIP image encoder are frozen all the time. During our training process, we find that dense sampling (30fps) results in jittering in generated videos. Therefore, we reduce the frame rate to one-fourth of the original, namely cutting the number of frames by three-quarters.
Optimization Objectives. We train Human2Robot with both generation and action prediction objectives, aiming for the model to achieve high-quality video generation and accurate action prediction simultaneously. The goal of generation for the model is to predict the applied noise, so the objective of the optimization process is:
@@
the parameters of the formula represent the latent state $z_{ts}$ at time step $ts$, which is obtained from Spatial-Temporal UNet. The conditioning variable $c$ includes the first frame $o^{r}_{t}$ of the robot and a human video $o^{h}_{t:t+n}$. $\epsilon$ denotes the added noise, and $\epsilon_{\theta}(z_{ts},c,ts)$ represents the model’s predicted noise.
@@ 4.3 Inference
Using KNN significantly simplifies the process for known tasks. However, for tasks involving different objectives within the same setting—such as pick cube task in Figure 4 and Table 2, as well as for unseen tasks such as handwriting, explicit human video guidance is required to accomplish these tasks effectively.
Temporal Addition of Actions. To mitigate trajectory jittering during sequential prediction, inspired by the approach used in ACT [72], we derive the action not from a single prediction but from a weighted average of multiple prediction steps. In particular, we predict a series of actions $a_{t:t+n}$ at time t, which are then combined with the previously predicted actions to compute the aggregated actions $A_{t:t+n}^{t}$. This represents the set of actions predicted at time $t$, to be executed from time $t$ to $t+n$, and can be computed as follows:
@@
where $\gamma$ is a hyperparameter, with a larger $\gamma$ indicating a stronger emphasis on the current prediction, and $m$ ($m\leq n$) refers to the length of the action executed after each prediction. Specifically, when $m=n$, the inference degenerates into sequential prediction.
Inference Acceleration. During inference, the diffusion model typically goes through $T$ denoising steps to generate the output. We explore the effectiveness of using features from different denoising steps in Section 5.4, and find that features obtained from the first half of the denoising stages lead to better results. Therefore, to balance speed and quality, we halt the denoising process after obtaining features from the $\frac{T}{4}$ step, directly predict and execute the action, without waiting for the entire denoising process to complete. It is worth noting that if we aim to achieve the fastest possible speed, we can directly use the one-step denoising features for prediction, which allows us to predict the next 30 steps of actions within 0.6 seconds.
Figure 4: Qualitative results. Visualization of generated result and robot action result.
@@ 5.1 Experimental Setups
Task Definition. As visualized in Figure 4, we conduct tests across the following four tasks:
  • Push the box. Push the box from left to right.
  • Pick the cube to the designated plate. Pick up the cube and place it on the bottom or top plate depending on the human demonstration.
  • Pick up and hold the brush upright. Pick up the brush, hold it upright, and get ready to write.
  • Handwriting. Write the corresponding character depending on the human demonstration.
The four carefully selected tasks above are designed to test the capabilities in multiple aspects: (1) The push box task is to test the accuracy in predicting height, position, and trajectory. (2) The pick cube task requires the model to understand the task objective in the video. (3) The upright task is a challenging fully 6-DOF task. (4) The handwriting task is to test the generalization ability.
Notably, in the handwriting task, the models are evaluated on unseen characters. Since there are countless characters in the world, it is impossible to demonstrate all of them. Therefore, having such a generalization ability is crucial. Thus, this task demands exceptional generalization capabilities, which we have not observed in other models.
Baselines. We compare with the following baselines:
  • RVCD Policy: We design a robot-video-conditioned policy $\pi(a_{t}|s_{t},v)$, where $v$ refers to the robot video that indicates the task. Specifically, we modify ACT [72] to a video conditioned policy. Details can be found in Supplementary materials D.2.
  • RVCD Policy w. Generated video: The framework and weights of this model is the same as RVCD Policy, but the robot video, generated by Human2Robot using human demonstration, is provided during inference.
  • XSkill [69]: XSkill is a human-video-conditioned policy tries to bridge the embodiment gap between human and robot. Details can be found in Supplementary materials D.2.
  • Human2Robot w. KNN Human2Robot with our KNN method proposed in Section 4.3, which enables to perform tasks without explicit demonstrations.
The baselines and Human2Robot utilize the same H&R dataset for training across all tasks except handwriting. For the handwriting task, the models are trained on handwriting play data [60] (randomly moving data).
Tasks Push Box Pick and Place Upright Handwriting
Apperence Position Instance Background Bg and Task
Generalization Type Brown Blue Pos1 Pos2 Pingpong Banana Bg1 Bg2 H R
XSkill [69]
Human2Robot
Table 1: Generalization evaluation. We evaluate generalization across four different tasks in unseen settings, including variations in appearances, positions, instances, backgrounds, and task targets. The scenes for each task are shown in the figure above. Detailed process can be found in the supplementary materials C. Each setting is evaluated with one trial.
@@ 5.2 Main results
Push Box Pick Cube Upright Handwriting
RVCD w. Gen 60 30 35 0
RVCD 75 45 40 0
XSkill [69] 70 40 50 0
Human2Robot w. KNN 85 40 60 0
Human2Robot 95 85 80 90
Table 2: Main results. Each task is evaluated with 20 trails.
Effectiveness. As shown in Table 2, Human2Robot achieves a high success rate across all tasks. In the pick cube to the designated plate task, half of the test cases require placing the cube on the bottom plate, while the other half require placing it on the top plate. Therefore, a success rate above 50% indicates that the model can correctly follow human demonstrations to perform distinct tasks. However, all baselines perform poorly on this task, consistently placing the cube on the bottom plate regardless of the given demonstration. This proves that only Human2Robot can truly understand human demonstration and execute the correct tasks. In addition, all baselines can successfully complete the handwriting task that requires strong generalization exhibitting they have limited generalization capabilities. These results highlight the satisfactory performance and strong generalization capabilities of Human2Robot.
We also test our KNN method proposed in Section 4.3, in a setting where no demonstrations are given. The performance of Human2Robot w. KNN is shown in Table 2. Human2Robot with KNN outperforms all other baselines across all tasks, demonstrating that even without direct demonstrations, Human2Robot can still achieve strong performance. However, the KNN method experiences a 10-20% drop in success rate on the push box and upright tasks, which is within an acceptable range. Meanwhile, a significant performance drop is observed in the pick cube to the designated plate task. This is expected, as the KNN method struggles to select the desired demonstration video when faced with the same initial scene. Finally, the KNN method cannot select demonstrations that were not present during training, as these demonstrations are only used during inference. As a result, it fails to accomplish the handwriting task. As a whole, such competitive results demonstrate the high efficiency and accuracy of Human2Robot in action prediction with episodes retrieved by KNN.
Quality of Generated Videos. Besides predicting actions, Human2Robot can visually simulate future dynamics of robot movements, requiring precise predictions of the motions of the robot arm and its interactions with objects. The visual quality of the generated video of Human2Robot is shown in Figure 4.
However, to evaluate the quality of generated videos, the traditional methods of using video distribution scores presents the following challenges: (1) Since the test set is small, which is common in the robotics domain, typically containing only 5 to 20 test samples, the scores may be unreliable. (2) No existing generation methods are directly applicable to our task.
To this end, we introduce a new visual quality evaluation pipeline: training a RVCD policy as an upper bound and replacing robot-conditioned videos with generated videos during inference. The performance drop serves as a proxy for evaluating visual quality. As shown in Table 2, the performance of RVCD w. Gen achieve a similar score, demonstrating the realism of our video prediction.
5.3 Generalization
Position generalization. Human2Robot uses its generative capability to enable the policy to understand human intentions and perform corresponding tasks at different positions. As shown in Table 1, Human2Robot achieves better results in different testing positions.
Appearance generalization. Appearance generalization is crucial for the model’s robustness. Since Human2Robot is based on Stable Diffusion [48], which contains a vast amount of prior knowledge, it naturally inherits strong generalization capabilities in terms of appearance without any data augmentation. As shown in the Table 1, Human2Robot demonstrates excellent generalization across different colors, which are unseen during training.
Instance generalization. Instance generalization requires the ability to generalize across objects of different sizes and shapes, making it even more critical than appearance generalization. As illustrated in the Table 1, thanks to the generative capability of diffusion models, Human2Robot achieves strong performance across various objects. Human2Robot successfully completes tasks on unseen instances, demonstrating its powerful generalization ability.
Background generalization. Although the training data primarily consists of white backgrounds, as shown in the Figure 4, we find that Human2Robot maintains stable performance even when the background color changes or the scene becomes cluttered. As presented in the Table 1, Human2Robot successfully completes tasks across different background, demonstrating its strong capability.
Task generalization. We train the model on play data, which involves moving randomly on the table rather than writing any specific character, and then evaluate its one-shot task execution ability given human handwriting prompts of never-before-seen character. To keep the scene clean during data collection, no actual handwriting was performed, resulting in blank backgrounds in the training data, which differ from those in the evaluation phase. Consequently, the handwriting task simultaneously challenges the generalization ability of models in both scene adaptation and task execution. The satisfying results shown in the Table 1 demonstrate the effectiveness and generalization capability of Human2Robot in task transfer.
@@ 5.4 Discussion
The Choice of Upsampling Layers for Feature Extraction. As shown in Table 3, we examine the impact of extracting features from different UNet upsample layers for action prediction. Using the 2nd up-sampling block (Up Index 2) achieves a 90% success rate on seen tasks, while the 1st up-sampling block (Up Index 1) reaches a 95% success rate on unseen tasks. The results indicate that the choice of upsampling layer has minimal effect on outcomes.
@@
Studying the Choice of Denoising Step. As shown in Tab 4, During inference, we fix the upsampling layer to extract features from different denoising steps. We find that, due to the MLP being trained with random noise, features from just one denoising step already allow the model to perform well, achieving 80% accuracy on seen tasks and 90% on unseen tasks, with an inference time of 0.6 seconds for 30 step trajectory prediction. Features after 7 denoising steps (1/4 of the total steps) yield the highest success rate. However, the results are worse at 15 (1/2 total steps) and 22 (3/4 total steps), possibly due to the following reason: the features in the early denoising stages, such as the 1st and 7th denoising stage shown in Figure 5, already capture sufficient spatial information to complete the tasks. However, in later denoising stages, the model focuses more on reconstructing detailed visual information, which increases action variance, making the movements jerky of the robot and ultimately reducing the success rate.
@@
Figure 5: Visual result of 1 and 7 step DDIM.
@@ 6 Conclusion
In this paper, we introduced a data collection pipeline for perfectly paired human-robot data and used it to create the H&R dataset. Then we presented Human2Robot, a novel end-to-end diffusion framework that simultaneously predicted robot video and actions given a hand video. Next, we leveraged a KNN method to perform tasks without explicitly requiring human hand videos as input. Human2Robot not only achieved outstanding performance on four carefully selected tasks but also demonstrated its capabilities in five generalization ability tests.